OpenStudy (anonymous):

What are the critical values of f(x)=8x^3-18x^2-240x+7

OpenStudy (anonymous):

poly. are continuous on (-infinity, infinity), so, find f'(x), set = 0, and solve.

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

24x^2-36x-240=0 2x^2-3x-20=0 x=-5/2, x=4

OpenStudy (anonymous):

you bet.